# DialectBench: A NLP Benchmark for Dialects, Varieties, and Closely-Related Languages

### v1

@@ Appendix F Highest performing and lowest performing varieties
DEP parsing (UD)
anglic/english*
italian romance/italian (r:formal, m:written, i:essay)*
tupi-guarani subgroup i.a/old guarani*
arabic/north african arabic
albanian/albanian*
@@
southwestern shifted romance/portuguese (a:european)
komi/komi-zyrian (m:written)*
italian romance/continental southern italian
gallo-rhaetian/french*
norwegian/norwegian nynorsk (m:written)*
saami/skolt saami*
komi/komi-zyrian (m:spoken)$\dagger$
norwegian/norwegian bokmål (m:written)*
southwestern shifted romance/portuguese (i:mix)*
tupi-guarani subgroup i.a/mbyá guaraní (a:paraguay)
komi/$\dagger$
italian romance/italian*
southwestern shifted romance/brazilian portuguese*
tupi-guarani subgroup i.a/mbyá guaraní (a:brazil)*
saami/north saami*$\dagger$
EQA (SD-QA-test)
anglic/english (a:scotland)
anglic/irish english
swahili/swahili (a:kenya)*
arabic/algerian $\dagger$
anglic/southern african english
anglic/philippine english
@@
bengali/vanga (a:west bengal)$\dagger$
arabic/tunisian $\dagger$
anglic/new zealand english
anglic/nigerian english
@@
bengali/vanga (a:dhaka)$\dagger$
arabic/moroccan $\dagger$
anglic/australian english
anglic/indian english (a:north)
korean/seoul (m:spoken)$\dagger$
arabic/egyptian arabic*$\dagger$
anglic/southeast american english*
anglic/kenyan english
@@
korean/korean (a:south-eastern, m:spoken)$\dagger$
swahili/swahili (a:tanzania)*
TC (SIB-200)
@@
sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$
italian romance/italian*
latvian/east latvian*
arabic/moroccan $\dagger$
anglic/english*
southwestern shifted romance/galician*
sotho-tswana (s.30)/northern sotho*
high german/limburgan
sinitic/cantonese*$\dagger$
southwestern shifted romance/spanish*
kurdish/northern $\dagger$
gallo-italian/lombard
@@
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
norwegian/norwegian nynorsk (m:written)*
sotho-tswana (s.30)/southern sotho*
gallo-italian/ligurian
southwestern shifted romance/portuguese (a:european)
arabic/standard arabic*$\dagger$
kurdish/central $\dagger$
common turkic/south $\dagger$
MRC (Belebele)
anglic/english*
arabic/north mesopotamian $\dagger$
arabic/moroccan $\dagger$
@@
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
@@
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
arabic/moroccan $\dagger$
arabic/egyptian arabic*$\dagger$
@@
sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$
sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$
arabic/egyptian arabic*$\dagger$
arabic/najdi $\dagger$
arabic/standard arabic*$\dagger$
arabic/standard arabic*$\dagger$
arabic/najdi $\dagger$
sotho-tswana (s.30)/northern sotho*
@@
arabic/levantine arabic (a:north)$\dagger$
@@
arabic/levantine arabic (a:north)$\dagger$
sotho-tswana (s.30)/northern sotho*
sotho-tswana (s.30)/southern sotho*
arabic/north mesopotamian $\dagger$
NER (Wikiann)
anglic/english (o:controlled)*
modern dutch/dutch*
sinitic/classical $\dagger$
mari/western $\dagger$
norwegian/norwegian nynorsk (m:written)*
southwestern shifted romance/galician*
sinitic/hakka chinese*$\dagger$
gallo-italian/emiliano-romagnolo
norwegian/norwegian (m:written, i:samnorsk)
italian romance/italian*
sotho-tswana (s.30)/northern sotho*
greater panjabic/eastern panjabi*$\dagger$
norwegian/norwegian bokmål (m:written)*
hindustani/fiji $\dagger$
kurdish/central $\dagger$
common turkic/south $\dagger$
anglic/english*
frisian/western frisian*$\dagger$
@@
anglic/jamaican creole english
kurdish/kurdish*$\dagger$
NLI (XNLI-translate-test)
anglic/english*
norwegian/norwegian bokmål (m:written)*
common turkic/south $\dagger$
arabic/north mesopotamian $\dagger$
southwestern shifted romance/spanish*
@@
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
kurdish/central $\dagger$
arabic/levantine arabic (a:north)$\dagger$
@@
southwestern shifted romance/portuguese (a:european)
southwestern shifted romance/occitan
sotho-tswana (s.30)/northern sotho*
high german/limburgan
italian romance/italian*
norwegian/norwegian nynorsk (m:written)*
sotho-tswana (s.30)/southern sotho*
arabic/tunisian $\dagger$
southwestern shifted romance/galician*
arabic/standard arabic*$\dagger$
kurdish/northern $\dagger$
latvian/east latvian*
POS tagging (UD)
anglic/english*
gallo-rhaetian/french (a:paris)
tupi-guarani subgroup i.a/mbyá guaraní (a:paraguay)
sinitic/classical $\dagger$
norwegian/norwegian bokmål (m:written)*
neva/finnish*
komi/komi-zyrian (m:written)*
arabic/levantine $\dagger$
high german/german*
italian romance/italian*
saami/skolt saami*
@@
italian romance/continental southern italian
norwegian/norwegian nynorsk (m:written)*
neva/estonian*
tupi-guarani subgroup i.a/old guarani*
komi/$\dagger$
gallo-rhaetian/french*
southwestern shifted romance/portuguese (a:european)
tupi-guarani subgroup i.a/mbyá guaraní (a:brazil)*
arabic/north african arabic
@@ Appendix H Cluster level Result Summaries with Demographic Utility and Standard Deviation report
albanian
63.3
69.1
(gheg albanian, 43.5) 28.0
anglic
75.5
90.7
(african american vernacular english, 52.5) 20.4
arabic
64.4
83.3
(south levantine arabic, 49.9) 21.0
eastern-western armenian
88.2
87.5
(eastern armenian, 87.0) 1.6
gallo-italian
50.2
50.2
(ligurian, 50.2) -
gallo-rhaetian
87.7
93.4
(french (a:paris), 77.5) 8.8
high german
56.6
76.5
(central alemannic (a:zh), 36.8) 28.1
italian romance
75.4
76.5
(continental southern italian, 50.0) 17.3
komi
30.2
30.1
(komi-zyrian (m:written), 27.6) 2.4
norwegian
88.4
(norwegian nynorsk (m:written, i:old), 78.4) 8.6
saami
47.9
67.0
(skolt saami, 28.3) 27.8
sabellic
33.2
(umbrian, 33.2) -
sinitic
65.4
(classical chinese, 46.7) 20.8
southwestern shifted romance
87.2
91.6
(portuguese (a:european), 77.3) 7.9
tupi-guarani subgroup i.a
17.2
22.0
(mbyá guaraní (a:brazil), 9.0) 10.5
west low german
40.8
40.8
(west low german, 40.8) -
@@
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
albanian
70.1 74.3
(gheg albanian, 55.8)
20.2
anglic
92.6 97.1
(singlish, 88.0)
6.5
arabic
61.9 81.4
(levantine arabic, 43.8)
18.9
eastern-western armenian
93.8 93.0
(eastern armenian, 92.4)
2.0
gallo-italian
58.9 58.9
(ligurian, 58.9)
gallo-rhaetian
96.4 96.9
(old french (842-ca. 1400), 95.6)
0.7
high german
75.5 88.4
(central alemannic (a:zh), 62.6)
18.2
italian romance
82.6 80.9
(continental southern italian, 57.1)
17.9
komi
41.8 41.6
(komi-zyrian (m:written), 35.1)
6.0
neva
75.4 97.7
(neva (a:south-west trans), 61.7)
13.9
norwegian
97.6 -
(norwegian nynorsk (m:written, i:old), 95.7)
1.6
saami
56.5 78.3
(skolt saami, 34.1)
31.6
sabellic
11.9 -
(umbrian, 11.9)
sinitic
84.5 -
(classical-middle-modern sinitic (a:hongkong, o:traditional), 69.0)
13.7
southwestern shifted romance
90.5 95.7
(occitan, 76.8)
10.5
tupi-guarani subgroup i.a
13.7 18.5
(mbyá guaraní (a:brazil), 1.9)
13.7
west low german
69.7 69.7
(west low german, 69.7)
@@
anglic
57.3 83.9 (jamaican creole english, 0.0) 40.9
arabic
81.2 86.0 (egyptian arabic, 73.3) 11.1
circassian
57.4 54.1 (kabardian, 47.5) 14.0
common turkic
67.9 85.1 (south azerbaijani, 31.7) 28.8
eastern romance
75.9 79.3 (moldavian, 64.6) 16.0
frisian
66.8 80.4 (northern frisian, 54.8) 13.3
gallo-italian
68.6 68.0 (emiliano-romagnolo, 45.5) 17.3
gallo-rhaetian
71.8 90.3 (walloon, 46.3) 16.5
greater panjabic
54.9 53.2 (eastern panjabi, 45.2) 13.6
greek
81.9 90.1 (pontic, 72.6) 13.2
high german
71.5 88.0 (pennsylvania german, 41.8) 15.9
hindustani
86.6 88.1 (fiji hindi, 85.2) 2.1
inuit
69.7 63.5 (kalaallisut, 63.2) 9.2
italian romance
73.0 88.4 (continental southern italian, 61.5) 13.6
komi
55.2 56.6 (komi-permyak, 53.6) 2.2
kurdish
47.5 54.8 (central kurdish, 35.5) 17.0
latvian
71.7 88.7 (east latvian, 50.1) 30.5
mari
52.6 47.4 (eastern mari, 46.7) 8.3
modern dutch
83.8 91.2 (zeeuws, 79.5) 7.1
norwegian
89.7 - (norwegian nynorsk (m:written), 87.9) 1.5
sardo-corsican
75.9 79.8 (corsican, 70.6) 7.5
serbian-croatian-bosnian
83.1 81.0 (serbian standard, 65.9) 11.9
sinitic
64.3 78.3 (hakka chinese, 40.0) 16.8
sorbian
68.3 67.4 (upper sorbian, 65.5) 3.9
sotho-tswana (s.30)
49.5 41.1 (northern sotho, 29.7) 28.0
southwestern shifted romance
74.0 92.1 (mirandese, 49.3) 18.1
west low german
80.3 80.3 (west low german, 80.3) -
@@
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
anglic
83.4 83.4
(english, 83.4)
arabic
63.3 70.0
(tunisian arabic, 50.2)
7.1
common turkic
64.1 62.3
(south azerbaijani, 44.6)
16.9
gallo-italian
61.7 63.6
(ligurian, 57.2)
6.0
gallo-rhaetian
54.6 54.6
(friulian, 54.6)
high german
59.9 59.8
(limburgan, 59.7)
0.2
italian romance
70.4 77.1
(sicilian, 62.7)
11.0
kurdish
51.4 55.5
(central kurdish, 39.6)
16.7
latvian
63.6 71.5
(east latvian, 53.5)
14.2
modern dutch
76.5 76.5
(dutch, 76.5)
norwegian
75.3 -
(norwegian nynorsk (m:written), 71.1)
6.0
sardo-corsican
58.3 58.3
(sardinian, 58.3)
sinitic
68.2 67.4
(classical-middle-modern sinitic (o:traditional), 64.5)
4.1
sotho-tswana (s.30)
35.3 35.6
(southern sotho, 34.6)
1.0
southwestern shifted romance
76.3 79.1
(occitan, 68.5)
5.2
@@
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
anglic 89.7 89.7
(english, 89.7)
arabic 84.5 85.7
(moroccan arabic, 79.1)
2.4
common turkic 78.7 77.3
(south azerbaijani, 69.7)
7.9
gallo-italian 73.8 73.7
(lombard, 70.6)
3.0
gallo-rhaetian 68.8 68.8
(friulian, 68.8)
high german 74.5 72.7
(limburgan, 71.1)
4.8
italian romance 81.4 86.8
(sicilian, 75.2)
8.8
kurdish 43.8 52.3
(central kurdish, 19.4)
34.5
latvian 75.6 82.0
(east latvian, 67.4)
11.5
modern dutch 89.6 89.6
(dutch, 89.6)
norwegian 86.7 -
(norwegian bokmål (m:written), 84.1)
3.6
sardo-corsican 71.0 71.0
(sardinian, 71.0)
sinitic 89.5 89.5
(classical-middle-modern sinitic (o:simplified), 89.2)
0.3
sotho-tswana (s.30) 37.8 36.9
(northern sotho, 35.6)
3.1
southwestern shifted romance 87.2 86.9
(occitan, 84.1)
2.3
@@
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
anglic 89.0 88.4
(north american english, 88.0)
1.4
arabic 58.1 89.0
(south levantine arabic (a:south-amm), 43.0)
11.3
greek 64.6 -
(cypriot greek (r:casual, m:written, i:twitter), 56.8)
6.8
high german 74.5 -
(central alemannic (a:be), 72.0)
2.1
sinitic 98.3 98.3
(mandarin chinese (a:mainland, o:traditional, i:synthetic), 97.9)
0.4
southwestern shifted romance 73.3 82.7
(portuguese (m:written), 17.4)
27.9
@@
clusters linguistic utility demographic utility
variety (minimum ling. u.)
standard deviation
anglic 53.4 53.4
(english, 53.4)
arabic 39.9 42.1
(moroccan arabic, 37.6)
2.4
sinitic 48.3 -
(classical-middle-modern sinitic (o:traditional), 46.9)
2.1
sotho-tswana (s.30) 30.1 30.5
(southern sotho, 29.0)
1.5

### v2

@@ Appendix F Highest performing and lowest performing varieties
DEP parsing (UD) anglic/english* italian romance/italian (r:formal, m:written, i:essay)* tupi-guarani subgroup i.a/old guarani* arabic/north african arabic
@@
albanian/albanian* southwestern shifted romance/portuguese (a:european) komi/komi-zyrian (m:written)* italian romance/continental southern italian
gallo-rhaetian/french* norwegian/norwegian nynorsk (m:written)* saami/skolt saami* komi/komi-zyrian (m:spoken)$\dagger$
norwegian/norwegian bokmål (m:written)* southwestern shifted romance/portuguese (i:mix)* tupi-guarani subgroup i.a/mbyá guaraní (a:paraguay) komi/$\dagger$
italian romance/italian* southwestern shifted romance/brazilian portuguese* tupi-guarani subgroup i.a/mbyá guaraní (a:brazil)* saami/north saami*$\dagger$
EQA (SD-QA-test) anglic/english (a:scotland) anglic/irish english swahili/swahili (a:kenya)* arabic/algerian $\dagger$
@@
anglic/southern african english anglic/philippine english bengali/vanga (a:west bengal)$\dagger$ arabic/tunisian $\dagger$
@@
anglic/new zealand english anglic/nigerian english bengali/vanga (a:dhaka)$\dagger$ arabic/moroccan $\dagger$
anglic/australian english anglic/indian english (a:north) korean/seoul (m:spoken)$\dagger$ arabic/egyptian arabic*$\dagger$
@@
anglic/southeast american english* anglic/kenyan english korean/korean (a:south-eastern, m:spoken)$\dagger$ swahili/swahili (a:tanzania)*
@@
TC (SIB-200) sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$ italian romance/italian* latvian/east latvian* arabic/moroccan $\dagger$
anglic/english* southwestern shifted romance/galician* sotho-tswana (s.30)/northern sotho* high german/limburgan
sinitic/cantonese*$\dagger$ southwestern shifted romance/spanish* kurdish/northern $\dagger$ gallo-italian/lombard
@@
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$ norwegian/norwegian nynorsk (m:written)* sotho-tswana (s.30)/southern sotho* gallo-italian/ligurian
southwestern shifted romance/portuguese (a:european) arabic/standard arabic*$\dagger$ kurdish/central $\dagger$ common turkic/south $\dagger$
@@
MRC (Belebele) anglic/english* arabic/north mesopotamian $\dagger$ arabic/moroccan $\dagger$ sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$
@@
sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$ arabic/moroccan $\dagger$ arabic/egyptian arabic*$\dagger$ sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$
@@
sinitic/classical-middle-modern sinitic (o:traditional)$\dagger$ arabic/egyptian arabic*$\dagger$ arabic/najdi $\dagger$ arabic/standard arabic*$\dagger$
@@
arabic/standard arabic*$\dagger$ arabic/najdi $\dagger$ sotho-tswana (s.30)/northern sotho* arabic/levantine arabic (a:north)$\dagger$
@@
arabic/levantine arabic (a:north)$\dagger$ sotho-tswana (s.30)/northern sotho* sotho-tswana (s.30)/southern sotho* arabic/north mesopotamian $\dagger$
NER (Wikiann) anglic/english (o:controlled)* modern dutch/dutch* sinitic/classical $\dagger$ mari/western $\dagger$
norwegian/norwegian nynorsk (m:written)* southwestern shifted romance/galician* sinitic/hakka chinese*$\dagger$ gallo-italian/emiliano-romagnolo
norwegian/norwegian (m:written, i:samnorsk) italian romance/italian* sotho-tswana (s.30)/northern sotho* greater panjabic/eastern panjabi*$\dagger$
norwegian/norwegian bokmål (m:written)* hindustani/fiji $\dagger$ kurdish/central $\dagger$ common turkic/south $\dagger$
@@
anglic/english* frisian/western frisian*$\dagger$ anglic/jamaican creole english kurdish/kurdish*$\dagger$
NLI (XNLI-translate-test) anglic/english* norwegian/norwegian bokmål (m:written)* common turkic/south $\dagger$ arabic/north mesopotamian $\dagger$
@@
southwestern shifted romance/spanish* sinitic/classical-middle-modern sinitic (o:simplified)$\dagger$ kurdish/central $\dagger$ arabic/levantine arabic (a:north)$\dagger$
@@
southwestern shifted romance/portuguese (a:european) southwestern shifted romance/occitan sotho-tswana (s.30)/northern sotho* high german/limburgan
italian romance/italian* norwegian/norwegian nynorsk (m:written)* sotho-tswana (s.30)/southern sotho* arabic/tunisian $\dagger$
southwestern shifted romance/galician* arabic/standard arabic*$\dagger$ kurdish/northern $\dagger$ latvian/east latvian*
POS tagging (UD) anglic/english* gallo-rhaetian/french (a:paris) tupi-guarani subgroup i.a/mbyá guaraní (a:paraguay) sinitic/classical $\dagger$
norwegian/norwegian bokmål (m:written)* neva/finnish* komi/komi-zyrian (m:written)* arabic/levantine $\dagger$
@@
high german/german* italian romance/italian* saami/skolt saami* italian romance/continental southern italian
norwegian/norwegian nynorsk (m:written)* neva/estonian* tupi-guarani subgroup i.a/old guarani* komi/$\dagger$
gallo-rhaetian/french* southwestern shifted romance/portuguese (a:european) tupi-guarani subgroup i.a/mbyá guaraní (a:brazil)* arabic/north african arabic
@@ Appendix H Cluster level Result Summaries with Demographic Utility and Standard Deviation report
albanian 63.3 69.1 (gheg albanian, 43.5) 28.0
anglic 75.5 90.7 (african american vernacular english, 52.5) 20.4
arabic 64.4 83.3 (south levantine arabic, 49.9) 21.0
eastern-western armenian 88.2 87.5 (eastern armenian, 87.0) 1.6
gallo-italian 50.2 50.2 (ligurian, 50.2) -
gallo-rhaetian 87.7 93.4 (french (a:paris), 77.5) 8.8
high german 56.6 76.5 (central alemannic (a:zh), 36.8) 28.1
italian romance 75.4 76.5 (continental southern italian, 50.0) 17.3
komi 30.2 30.1 (komi-zyrian (m:written), 27.6) 2.4
norwegian 88.4 - (norwegian nynorsk (m:written, i:old), 78.4) 8.6
saami 47.9 67.0 (skolt saami, 28.3) 27.8
sabellic 33.2 - (umbrian, 33.2) -
sinitic 65.4 - (classical chinese, 46.7) 20.8
southwestern shifted romance 87.2 91.6 (portuguese (a:european), 77.3) 7.9
tupi-guarani subgroup i.a 17.2 22.0 (mbyá guaraní (a:brazil), 9.0) 10.5
west low german 40.8 40.8 (west low german, 40.8) -
@@
albanian 70.1 74.3 (gheg albanian, 55.8) 20.2
anglic 92.6 97.1 (singlish, 88.0) 6.5
arabic 61.9 81.4 (levantine arabic, 43.8) 18.9
eastern-western armenian 93.8 93.0 (eastern armenian, 92.4) 2.0
gallo-italian 58.9 58.9 (ligurian, 58.9) -
gallo-rhaetian 96.4 96.9 (old french (842-ca. 1400), 95.6) 0.7
high german 75.5 88.4 (central alemannic (a:zh), 62.6) 18.2
italian romance 82.6 80.9 (continental southern italian, 57.1) 17.9
komi 41.8 41.6 (komi-zyrian (m:written), 35.1) 6.0
neva 75.4 97.7 (neva (a:south-west trans), 61.7) 13.9
norwegian 97.6 - (norwegian nynorsk (m:written, i:old), 95.7) 1.6
saami 56.5 78.3 (skolt saami, 34.1) 31.6
sabellic 11.9 - (umbrian, 11.9) -
sinitic 84.5 - (classical-middle-modern sinitic (a:hongkong, o:traditional), 69.0) 13.7
southwestern shifted romance 90.5 95.7 (occitan, 76.8) 10.5
tupi-guarani subgroup i.a 13.7 18.5 (mbyá guaraní (a:brazil), 1.9) 13.7
west low german 69.7 69.7 (west low german, 69.7) -
@@
anglic 57.3 83.9 (jamaican creole english, 0.0) 40.9
arabic 81.2 86.0 (egyptian arabic, 73.3) 11.1
circassian 57.4 54.1 (kabardian, 47.5) 14.0
common turkic 67.9 85.1 (south azerbaijani, 31.7) 28.8
eastern romance 75.9 79.3 (moldavian, 64.6) 16.0
frisian 66.8 80.4 (northern frisian, 54.8) 13.3
gallo-italian 68.6 68.0 (emiliano-romagnolo, 45.5) 17.3
gallo-rhaetian 71.8 90.3 (walloon, 46.3) 16.5
greater panjabic 54.9 53.2 (eastern panjabi, 45.2) 13.6
greek 81.9 90.1 (pontic, 72.6) 13.2
high german 71.5 88.0 (pennsylvania german, 41.8) 15.9
hindustani 86.6 88.1 (fiji hindi, 85.2) 2.1
inuit 69.7 63.5 (kalaallisut, 63.2) 9.2
italian romance 73.0 88.4 (continental southern italian, 61.5) 13.6
komi 55.2 56.6 (komi-permyak, 53.6) 2.2
kurdish 47.5 54.8 (central kurdish, 35.5) 17.0
latvian 71.7 88.7 (east latvian, 50.1) 30.5
mari 52.6 47.4 (eastern mari, 46.7) 8.3
modern dutch 83.8 91.2 (zeeuws, 79.5) 7.1
norwegian 89.7 - (norwegian nynorsk (m:written), 87.9) 1.5
sardo-corsican 75.9 79.8 (corsican, 70.6) 7.5
serbian-croatian-bosnian 83.1 81.0 (serbian standard, 65.9) 11.9
sinitic 64.3 78.3 (hakka chinese, 40.0) 16.8
sorbian 68.3 67.4 (upper sorbian, 65.5) 3.9
sotho-tswana (s.30) 49.5 41.1 (northern sotho, 29.7) 28.0
southwestern shifted romance 74.0 92.1 (mirandese, 49.3) 18.1
west low german 80.3 80.3 (west low german, 80.3) -
@@
anglic 83.4 83.4 (english, 83.4) -
arabic 63.3 70.0 (tunisian arabic, 50.2) 7.1
common turkic 64.1 62.3 (south azerbaijani, 44.6) 16.9
gallo-italian 61.7 63.6 (ligurian, 57.2) 6.0
gallo-rhaetian 54.6 54.6 (friulian, 54.6) -
high german 59.9 59.8 (limburgan, 59.7) 0.2
italian romance 70.4 77.1 (sicilian, 62.7) 11.0
kurdish 51.4 55.5 (central kurdish, 39.6) 16.7
latvian 63.6 71.5 (east latvian, 53.5) 14.2
modern dutch 76.5 76.5 (dutch, 76.5) -
norwegian 75.3 - (norwegian nynorsk (m:written), 71.1) 6.0
sardo-corsican 58.3 58.3 (sardinian, 58.3) -
sinitic 68.2 67.4 (classical-middle-modern sinitic (o:traditional), 64.5) 4.1
sotho-tswana (s.30) 35.3 35.6 (southern sotho, 34.6) 1.0
southwestern shifted romance 76.3 79.1 (occitan, 68.5) 5.2
@@
anglic 89.7 89.7 (english, 89.7) -
arabic 84.5 85.7 (moroccan arabic, 79.1) 2.4
common turkic 78.7 77.3 (south azerbaijani, 69.7) 7.9
gallo-italian 73.8 73.7 (lombard, 70.6) 3.0
gallo-rhaetian 68.8 68.8 (friulian, 68.8) -
high german 74.5 72.7 (limburgan, 71.1) 4.8
italian romance 81.4 86.8 (sicilian, 75.2) 8.8
kurdish 43.8 52.3 (central kurdish, 19.4) 34.5
latvian 75.6 82.0 (east latvian, 67.4) 11.5
modern dutch 89.6 89.6 (dutch, 89.6) -
norwegian 86.7 - (norwegian bokmål (m:written), 84.1) 3.6
sardo-corsican 71.0 71.0 (sardinian, 71.0) -
sinitic 89.5 89.5 (classical-middle-modern sinitic (o:simplified), 89.2) 0.3
sotho-tswana (s.30) 37.8 36.9 (northern sotho, 35.6) 3.1
southwestern shifted romance 87.2 86.9 (occitan, 84.1) 2.3
@@
anglic 89.0 88.4 (north american english, 88.0) 1.4
arabic 58.1 89.0 (south levantine arabic (a:south-amm), 43.0) 11.3
greek 64.6 - (cypriot greek (r:casual, m:written, i:twitter), 56.8) 6.8
high german 74.5 - (central alemannic (a:be), 72.0) 2.1
sinitic 98.3 98.3 (mandarin chinese (a:mainland, o:traditional, i:synthetic), 97.9) 0.4
southwestern shifted romance 73.3 82.7 (portuguese (m:written), 17.4) 27.9
@@
anglic 53.4 53.4 (english, 53.4) -
arabic 39.9 42.1 (moroccan arabic, 37.6) 2.4
sinitic 48.3 - (classical-middle-modern sinitic (o:traditional), 46.9) 2.1
sotho-tswana (s.30) 30.1 30.5 (southern sotho, 29.0) 1.5